OpenStudy (anonymous):

What is the fraction 36/80 expressed in lowest terms?

OpenStudy (anonymous):

factors of 36 divided by factors of 80 cancel the common factors

OpenStudy (anonymous):

Divide both by 4

OpenStudy (anonymous):

why dont you divide them by 2\

OpenStudy (anonymous):

36/2 80/2

OpenStudy (anonymous):

36/80 = 18/40 keep simplifying it.

OpenStudy (anonymous):

9/20

OpenStudy (anonymous):

\(\dfrac{36\ \div\ 4}{80\ \div\ 4}\)

OpenStudy (anonymous):

@TreyYoung Please do not give direct answers. Since you are new here, you can read the rules here: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@AnswerMyQuestions would be right

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

@casbywasby Can you divide 36 \(\div\) 4 and 80 \(\div\) 4?